If the moment generating function (m.g.f.) of a random variable X is 

Mx(t)=exp(3t+32t2).Find mean and standard derivation of X and also compute 

P(X<3).


1
Expert's answer
2021-06-03T09:26:35-0400

Solution:

Given, "M_X(t)=e^{(3t+32t^2)}" ...(i)

Now, mean"=\mu=M'_X(0)"

On differentiating (i) w.r.t t,

"M'_X(t)=e^{(3t+32t^2)}(3+64t)" ...(ii)

Put "t=0"

"M'_X(0)=e^{0}(3)=3"

Thus, mean = 3

Again, differentiating (ii) w.r.t. t,

"M''_X(t)=[e^{(3t+32t^2)}]'(3+64t)+e^{(3t+32t^2)}(3+64t)'
\\=e^{(3t+32t^2)}(3+64t)^2+e^{(3t+32t^2)}(64)
\\=e^{(3t+32t^2)}[64+(3+64t)^2]"

Put "t=0"

"M''_X(0)=e^{0}[64+(3)^2]=1(73)=73"

Now, "\sigma^2=M''_X(0)-[M'_X(0)]^2=73-3^2=64"

Then, "\sigma=8"

Assume "X\sim N(\mu,\sigma)"

"P(X<3)=P(z<\dfrac{3-3}{8})=P(z<0)=0.5"